Homework 3 11-721: Grammars and Lexicons, Fall 2007 Due Monday, October 8, 2007

1 Kroeger, Exercise 3B, Pages 77-79

Using the data in Kroeger’s Exercise 3B, follow these instructions:

Task 1: In Sentence B1, the arguments of the verb are the subject, guru kami (our teacher), and the oblique di ruang Pak rektor (in the rector’s office). What are the arguments of the verb in Sentence B2?

Task 2: Look at the verbs of Sentences B1 and B2. (a) Aside from the prefix men-, which is not relevant to this exercise, what affix is added to the verb in sentence B2? (b) What effect does the affix have on the sentence?

Task 3: Sentence B3 is a passive sentence. (a) Which morpheme is added to the verb to make it passive? (b) Assuming that the subject comes to the left of the verb in Indonesian, what is the subject of the sentence?

Note: You cannot rely on the English translation to tell you the grammatical relations (subject, object, oblique) in Indonesian. English does not have applicative morpheme, so the grammatical relations are different in English.

Task 4: Finish the following lexical entry for the verb in Sentence B1 by adding a grammatical relation for each semantic role:

duduk <  theme    location  >

Task 5: Finish the lexical entry for the verb in Sentence B2 by adding a grammatical relation for each semantic role:

duduk <  theme    location  >

Task 6: Finish the lexical entry for the verb in Sentence B3 by adding a grammatical relation for each semantic role:

duduk <  theme    location  >

Task 7: List three more sentences that illustrate the same phenomena as B1, B2, and B3.

Task 8: Consider two hypotheses:

• Hypothesis 1: Sentence B4 has a subject and an object.

• Hypothesis 2: Sentence B4 has a subject and an oblique.

Assume the following passive rule: Find the semantic role that was linked to the grammatical relation object in the active sentence and link it to the grammatical relation subject. Find the semantic role that was linked to subject in the active sentence and link it to oblique.

Sentence B7 is ungrammatical. Which hypothesis does it support? Why?

Task 9: Compare Sentences B12 and B13. (a) What suffix is added to the verb in B13? (b) What does it do?

Task 10: Complete the lexical entry for Sentence B12:

beli <  agent  theme  beneficiary  >

Task 11: Complete the lexical entry for Sentence B13:

beli <  agent  theme  beneficiary  >

Task 12: (a) Complete the lexical entry for Sentence B14. (b) Is sentence B14 the passive of Sentence B12 or B13?

beli <  agent  theme  beneficiary  >

Task 13: (a) Complete the lexical entry for Sentence B15. (b) Is sentence B15 the passive of Sentence B12 or B13?

beli <  agent  theme  beneficiary  >

Task 14: Why is Sentence B16 ungrammatical?

Task 15: Identify four more sentences that illustrate the same points as B12-B15.

2 Kinyarwanda

Van Valin and Kroeger both discuss data from Kinyarwanda.

Task 1: List two sentences in Kroeger’s Exercise 3B that are ungrammatical in Indonesian, but probably would be grammatical in Kinyarwanda, based on Van Valin’s description of Kinyarwanda (pages 64 and 65). The answer has something to do with whether the theme or the applicative NP passes tests for objecthood. I’m looking for a very specific and short answer.

Task 2: Kinyarwanda has more applicative suffixes than Chichewa. Based on Kroeger’s description of Kinyarwanda, Section 3.3.2 identify a sentence in Kinyarwanda that is probably not possible in Chichewa. The answer will be based on the semantic roles that are affected by the applicative suffixes in the two languages.

3 Extra Credit

Look at Keenan, Passive in the world’s languages. List facts that you didn’t know before or that you found surprising, or that you have questions about. Include a page number for each thing on your list. You will get one quarter percent extra credit for each correct item or reasonable question that you list. You can list as many items and questions as you want, but I will only count them if I feel that you understand what you are commenting on. (Don’t say, “I didn’t know that X” where X is copied word-for-word from the paper.)